Consider a 1.0 L buffer containing 0.092 mol L-1 HCOOH and 0.100 mol L-1 HCOO-. What is the pH of the solution after adding 1.9 x 10-3 mol of HCl? Express your answer to 2 decimal places.
Answer #1

Molarity of hcl = moles / volume

= 1.9*10^-3 / 1

= 1.9*10^-3 M

Now it will reduce the amount of conjugate base and increases the amount of acid

Base = 0.1 - (1.9*10^-3 )

= 0.0981

Acid = 0.092 + (1.9*10^-3 )

= 0.0939

pKa of HCOOH = 3.8

pH = pKa + log ( conjugate base / acid )

= 3.8 + log ( 0.0981 / 0.0939)

= 3.82

If you have any query please comment

If you satisfied with the solution please like it thankxx